# Learning to Evolve with Convergence Guarantee via Neural Unrolling

### v1

@@ Abstract
The transition from hand-crafted heuristics to data-driven evolutionary algorithms faces a fundamental dilemma: achieving neural plasticity without sacrificing mathematical stability. Emerging learned optimizers demonstrate high adaptability. However, they often lack rigorous convergence guarantees. This deficiency results in unpredictable behaviors on unseen landscapes. To address this challenge, we introduce Learning to Evolve (L2E), a unified bilevel meta-optimization framework. This method reformulates evolutionary search as a Neural Unrolling process grounded in Krasnosel’skii-Mann (KM) fixed-point theory. First, L2E models a coupled dynamic system in which the inner loop enforces a strict contractive trajectory via a structured Mamba-based neural operator. Second, the outer loop optimizes meta-parameters to align the fixed point of the operator with the target objective minimizers. Third, we design a gradient-derived composite solver that adaptively fuses learned evolutionary proposals with proxy gradient steps, thereby harmonizing global exploration with local refinement. Crucially, this formulation provides the learned optimizer with provable convergence guarantees. Extensive experiments demonstrate the scalability of L2E in high-dimensional spaces and its robust zero-shot generalization across synthetic and real-world control tasks. These results confirm that the framework learns a generic optimization manifold that extends beyond specific training distributions.
@@ I Introduction
Evolutionary Algorithms (EAs) constitute a core paradigm in computational intelligence for solving complex black-box optimization problems [1, 2, 3, 4, 5]. These algorithms evolve a population of candidate solutions through selection, recombination, and perturbation. This mechanism enables them to explore highly nonconvex, multimodal, and noisy landscapes without relying on explicit gradient information. Consequently, this paradigm plays a foundational role in computational science and engineering. It underpins diverse applications such as hyperparameter tuning in machine learning [6], aerodynamic shape design [7], molecular conformation prediction [8], and the control of physical systems [9]. Specifically, solvers in these scenarios must navigate high-dimensional and noisy search spaces under tight evaluation budgets. They must simultaneously achieve effective exploration and reliable convergence to high-quality solutions [10].
Optimization research has established two principal families over the past decades: gradient-free, population-based Evolutionary Algorithms (EAs) and model-based Bayesian Optimization (BO). The former family, which includes Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [11], Differential Evolution (DE) [12], and various swarm-based methods [13], is noted for conceptual simplicity and global exploration. Conversely, the latter family is widely adopted for sample efficiency in computationally expensive contexts [14]. However, both families rely on hand-crafted operators and heuristics. Furthermore, they lack a unified theoretical foundation for convergence and stability. This inherent limitation constrains their efficiency in high-dimensional spaces. Consequently, this results in a fundamental lack of adaptability to increasingly complex modern workloads.
A natural response to these limitations is the shift from manually designing operators to data-driven and learnable evolutionary optimizers [15, 16, 17]. Within the Learning to Optimize (L2O) paradigm, recent studies leverage Recurrent Neural Networks (RNNs) and Transformers as controllers, or employ set-based architectures to model optimization trajectories [18]. In parallel, meta-evolution approaches treat mutation-crossover-selection policies as learnable objects. These methods search over expressive families of attention-based, diffusion-inspired, or transformer-style operators [19, 20, 21]. Specifically, they seek to automatically discover evolutionary strategies by learning from a task distribution, such as LQD [22] and the Evolution Transformer [23].
However, a critical gap remains in this transition: the emerging learned optimizers typically face a fundamental dilemma between plasticity and stability. Most existing approaches treat the optimizer as a black-box policy network trained via reinforcement learning to simply minimize trajectory loss. While they demonstrate exceptional plasticity on training tasks, they frequently degenerate into mathematically unconstrained processes. Lacking intrinsic geometric constraints, they are prone to overfitting the training landscape topology, figuratively “memorizing the map” rather than “learning the compass”. Consequently, when deployed on unseen or out-of-distribution tasks, these solvers often exhibit wandering trajectories, catastrophic divergence, or stagnation, failing to provide the stability and rigorous convergence assurances characterizing classical EAs.
Consequently, this limitation motivates us to rethink the fundamental design of learned optimizers: how can we engineer a learned evolutionary optimizer that possesses the adaptive intelligence of neural networks without sacrificing the rigorous convergence laws of classical control? To address this challenge, we propose to ground the design of learned EAs in fixed-point theory. We argue that the fragility of current learned optimizers stems from unconstrained operator learning. A robust evolutionary operator should not merely predict the next step based on heuristics; rather, it must mathematically constitute a contraction mapping. The repeated application of this mapping drives the population towards equilibrium.
Motivated by this insight, we introduce Learning to Evolve (L2E), a unified framework that reformulates evolutionary search as a neural unrolling process. Unlike approaches that view the optimizer merely as a meta-learner, L2E models the optimizer as a learnable dynamical system governed by the classical Krasnosel’skiĭ-Mann (KM) iteration [24]. The core philosophy is to enforce a physics of descent within the neural architecture. In the inner loop, the population update serves as an $\alpha$-averaged operator composed of a learned component and a numerical proxy. This structure enforces non-expansiveness by design. Consequently, the evolution trajectory strictly contracts towards a fixed point, ensuring stability regardless of the landscape’s complexity. In the outer loop, the meta-learner aligns the operator’s fixed point with the target objective minimizers. This alignment effectively shapes the attraction basin of the dynamics. To implement this efficiently in high-dimensional spaces, we propose a structured Mamba-based operator, where spectral normalization constraints are imposed. By leveraging the linear complexity of State Space Models (SSMs), this design overcomes the quadratic bottleneck of Transformers in population-wise interaction complexity, capturing long-range dependencies dynamics efficiently.
This architectural formulation yields the behavioral contrast illustrated in Fig. 1. While conventional learned optimizers often succumb to overfitting the training landscape, which results in unpredictable or unstable trajectories on unseen tasks, L2E induces a stable contraction field where the learned fixed point is explicitly aligned with the target optimum. Consequently, the proposed L2E avoids the pitfall of memorizing the map by learning a generalized descent rule, thereby offering a robust solution for general-purpose adaptation. Our main contributions are:
  • Theory-Driven Formulation: We introduce the Generalized Neural Unrolling (GNU) paradigm. This framework redefines evolution as a curriculum-driven unrolling process under bilevel fixed-point constraints. Furthermore, we establish a joint convergence guarantee to bridge neural plasticity with theoretical optimization rigor.
  • Efficient Composite Solver: We design a Gradient-derived Evolutionary Composite (GEC) solver, which utilizes a Mamba architecture and an elite-mask fusion mechanism. This design addresses the scalability limitations of attention-based optimizers. Consequently, it enables stable meta-evolution in high-dimensional regimes where traditional L2O methods struggle.
  • Rigorous Theoretical Analysis: We provide a fixed-point theoretic analysis establishing the finite-time convergence rate, asymptotic consistency, and bilevel optimality. These guarantees ensure the learned optimizer follows a predictable trajectory, avoiding the instability inherent in unconstrained meta-learning.
@@
Figure 1: Conceptual contrast. Left: Naive optimizers memorize the map, leading to unstable wandering on unseen landscapes. Right: L2E enforces a physics of descent via non-expansive unrolling, inducing a stable contraction field where the learned fixed-point is aligned with the target optimum.
@@ II-A Learning to Optimize
Learning to Optimize (L2O) blends the prescriptive rigor of classical optimization with the inductive power of machine learning [18]. Specifically, rather than relying on hand-crafted heuristics, L2O seeks to parameterize update rules directly from data. Early works employed model-free architectures, such as RNNs [25] or MLPs [26], to map gradients to updates in a coordinate-wise manner. However, these methods often struggle with scalability and interpretability, primarily due to their opaque nature.
In contrast, model-based L2O approaches embed structural priors into the learner. For instance, Transformer-based optimizers [27] leverage self-attention to capture parameter dependencies. This methodology has further evolved into meta-learning frameworks that train optimizers to adapt across diverse tasks [20, 28]. Notable implementations include reinforcement learning controllers for scheduling [19] and gradient-guided mutation operators [23].
Despite these advances, existing L2O approaches primarily function as unconstrained policy networks. Specifically, they lack intrinsic geometric constraints. This absence implies that there is no theoretical assurance of convergence or stability. Consequently, this deficiency frequently leads to unpredictable behavior when the model transfers to out-of-distribution landscapes, thereby highlighting the necessity for mathematically grounded architectures that enforce stability by design.
@@ II-B Learnable Evolutionary Algorithms
Classical evolutionary algorithms and Bayesian Optimization methods have established the foundation for gradient-free optimization. Prominent examples include Evolutionary Strategies, such as CMA-ES [11], Differential Evolution [12], and L-SHADE [13]. Within the Bayesian domain, notable approaches include SAASBO [29], TuRBO [30], and HEBO [14]. These methods proved successful due to engineered covariance adaptation and population mechanisms [31]. However, their reliance on hand-crafted, non-differentiable rules limits their flexibility in high-dimensional or ill-conditioned landscapes.
To enhance adaptability, recent research has progressively integrated learnability into evolutionary frameworks [32, 33, 34]. Initial efforts utilized surrogates, such as Gaussian Processes, to approximate objective landscapes [35]. Subsequent work has focused on embedding data-driven modules directly into the optimization loop. Examples include meta-evolved operators [36] and context-aware optimizers [19, 20]. Concurrently, researchers have employed Reinforcement Learning for dynamic operator selection [37, 38, 39]. Furthermore, Large Language Models are increasingly being explored for synthesizing heuristic code [40, 41, 42].
Despite this progress, learnable EAs face substantial challenges. Most existing works optimize isolated components rather than the unified evolutionary loop. This isolated optimization results in a fragmented learning process. Furthermore, current methods often lack a consistent mathematical framework to integrate gradient propagation with population search. Consequently, achieving robust zero-shot generalization and scalability in high-dimensional regimes remains a persistent difficulty.
@@ II-C Neural Unrolling
Within the broader inverse problem community, Neural Unrolling (NU) has established itself as a principled methodology for bridging model-based iteration and deep learning. Fundamentally, NU reinterprets an iterative optimization method as a computational graph, wherein each layer corresponds to a single update step and algorithmic parameters function as trainable weights [43]. Although originally developed for signal and image processing, this paradigm has produced influential architectures, including LISTA for sparse coding [44], ADMM-Net for compressive sensing [45], and primal-dual networks for medical image reconstruction [46]. NU embeds domain-specific iterations into differentiable architectures. Recent advances have extended NU to complex visual perception tasks, such as image dehazing and rain removal, by incorporating optimization-inspired heuristics into network design to enhance feature interpretability [47, 48, 49]. Furthermore, in the realm of AutoML, differentiable approaches for Neural Architecture Search (NAS) implicitly leverage the unrolling concept to optimize structural parameters via bilevel gradients [50, 51, 52]. Consequently, it synergizes the interpretability of model-based methods with the adaptability of deep networks, thereby achieving robust performance across diverse inverse problems.
In contrast to its prevalence in signal processing, NU remains underutilized within evolutionary computation. While differentiable EAs exist, they typically rely on ad hoc relaxations rather than systematic unrollings of fixed-point iterations. Thus, the field lacks a framework capable of unifying bilevel meta-learning with rigorous population dynamics guarantees.
@@ II-D Discussion
Existing methodologies navigate a critical trade-off between stability and adaptability. While classical EAs provide theoretical stability, they often lack the adaptability requisite for navigating complex landscapes efficiently. Conversely, emerging L2O approaches offer data-driven flexibility; however, they frequently exhibit instability and poor generalization to out-of-distribution landscapes, primarily due to the absence of intrinsic geometric constraints. To reconcile these diverging paradigms, we propose the L2E that reformulates the evolutionary loop as a neural unrolling process. This formulation is expected to synergize the expressivity of deep networks with the rigorous convergence properties of classical optimization.
Figure 2: Overall framework of the proposed L2E paradigm. Left: Visualizes the interaction where the outer-loop meta-learner reshapes the landscape to align the inner-loop fixed point with the target optimum. Middle: Illustrates the solving process, where the evolutionary search is unrolled into a differentiable pipeline to learn stable trajectories. Right: Details the non-expansive unrolling operator integrates evolutionary and numerical solvers to form a differentiable evolution dynamics.
@@ III Methodology
In this section, we present L2E, a framework designed to unify heuristic search and mathematical optimization. The overall architecture functions as a bilevel dynamical system, as visualized in Fig. 2. To instantiate this paradigm, we first introduce the Generalized Neural Unrolling (GNU) formulation, which provides the theoretical backbone. Second, we detail the Gradient-derived Evolutionary Composite (GEC) solver, which harmonizes gradient-based exploitation with population-based exploration. Finally, we propose a Structured Mamba Neural Operator to efficiently parameterize this process in high-dimensional spaces, thereby ensuring both computational scalability and theoretical stability.
@@ III-A Generalized Neural Unrolling Paradigm
Conventional black-box optimizers often rely on heuristic mechanisms that lack theoretical guarantees. To address this limitation, we reinterpret the optimization process through the lens of non-expansive operator theory. Specifically, we utilize the KM fixed-point formulation [24]. Unlike methods that directly search for optima, the proposed approach optimizes the parameters of a learnable fixed-point operator. This operator implicitly governs the solution trajectory. Formally, let $f:\mathcal{X}\to\mathbb{R}$ denote the black-box objective function. We introduce a learnable non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$, parameterized by $\bm{\omega}\in\bm{\Omega}$, whose fixed point approximates the minimizer of $f$. Consequently, we formulate this framework as a constrained bilevel optimization problem [53]:
@@
where $\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt})$ denotes a meta-level surrogate objective that encodes task rewards, diversity, or other downstream fitness metrics. Furthermore, $\text{Fix}(\mathcal{NU}):=\{\bm{\mathbf{x}}\mid\bm{\mathbf{x}}=\mathcal{NU}(\bm{\mathbf{x}};\bm{\omega})\}$ defines the set of fixed points. In this formulation, $\bm{\omega}$ represents the outer-level parameters defining the optimization algorithm, whereas $\bm{\mathbf{x}}$ denotes the inner-level population of candidate solutions.
To ensure contractive behavior, we construct $\mathcal{NU}$ as an $\alpha$-averaged operator. Specifically, this formulation employs a convex combination of the identity and a learnable mapping $\mathcal{O}(\cdot,\bm{\omega})$, thereby yielding a KM-style operator:
@@
The operator $\mathcal{NU}(\cdot;\bm{\omega})$ is defined as a composition of two components, modulated by an update relaxation parameter $\alpha\in(0,1]$. Specifically, $\mathcal{O}_{\text{num}}$ denotes a numerical solver, such as smoothing or projection, while $\mathcal{O}_{\text{evo}}$ encapsulates a learnable search policy. By enforcing non-expansiveness on $\mathcal{O}$, the GNU paradigm ensures that the evolution trajectory $\mathbf{x}^{0}\to\dots\to\mathbf{x}^{K}$ follows a strictly contractive path. Consequently, this formulation provides a theoretical safeguard against divergence, which addresses a common failure mode in standard learned optimizers.

### v2

# Learning to Evolve for Optimization via Stability-Inducing Neural Unrolling

@@ Abstract
Evolutionary algorithms serve as a powerful paradigm for tackling optimization challenges, yet their reliance on manually engineered heuristics inherently limits their adaptability across diverse landscapes. However, the transition from the hand-crafted heuristics to data-driven algorithms faces a fundamental dilemma: achieving neural plasticity without sacrificing algorithmic stability. Although learned optimizers offer high adaptivity, their unconstrained update rules often result in unstable dynamics and brittle generalization on unseen landscapes. To address this challenge, this paper proposes Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search via stability-inducing neural unrolling. First, L2E reformulates population evolution as an unrolled fixed-point iteration via a structured neural operator. In this design, the inner loop imposes a stability-biased update structure, while the outer loop meta-trains the operator to produce effective search trajectories across tasks. Second, to balance global exploration with local refinement, a gradient-derived composite solver adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner. Extensive experiments on synthetic benchmarks and real-world control tasks demonstrate that L2E achieves substantial optimization performance, scales to high-dimensional problems, and exhibits robust zero-shot transfer across diverse test distributions.
@@ I Introduction
Evolutionary Algorithms (EAs) are a core paradigm in computational intelligence for solving complex black-box optimization problems [67, 46, 65, 8, 5]. By evolving a population through selection, recombination, and perturbation, EAs can explore highly nonconvex, multimodal, and noisy landscapes without relying on explicit gradient information. As a result, they underpin a wide range of applications, including hyperparameter tuning [14], aerodynamic shape design [24], molecular conformation prediction [51], and the control of physical systems [15]. In these scenarios, solvers must operate under tight evaluation budgets while navigating high-dimensional and noisy search spaces, requiring both effective exploration and reliable progress toward high-quality solutions [9].
Over the past decades, black-box optimization has been dominated by two major families: population-based evolutionary search and model-based Bayesian optimization (BO). The former emphasizes exploration through handcrafted operators and heuristics, whereas the latter focuses on sample efficiency via surrogate modeling in expensive evaluation settings [12]. Despite their empirical success, a common bottleneck is that the update dynamics are largely driven by manually designed heuristics. Consequently, their behavior is difficult to characterize and control in a unified way across varying scales, conditioning, and dimensionalities. This issue becomes particularly pronounced in high-dimensional regimes and when the optimizer is deployed on unseen tasks, where stable progress under limited evaluations is challenging.
A natural response is the transition from manually designing operators to data-driven and learnable evolutionary optimizers [62, 70, 66, 63]. Recent studies in learning to optimize (or meta-evolution) learn mutation-crossover-selection policies from a task distribution. Specifically, these methods employ neural controllers (e.g., RNNs, Transformers) or set-based architectures to model population updates [6, 3, 44, 45, 42]. Representative examples include learning-driven operator discovery and attention-based evolutionary strategies, such as LQD [4] and the evolution transformer [32]. However, a critical gap remains: learned optimizers often struggle to reconcile plasticity with stability in a way that transfers reliably across tasks. Most existing approaches learn update rules as unconstrained policies that minimize trajectory loss on training problems. While such policies can be highly effective in-distribution, they may induce unstable dynamics on unseen landscapes, such as oscillation, premature stagnation, or sensitivity to nuisance shifts in scale and conditioning. In other words, the learned update rule may implicitly overfit the training landscape topology, i.e., “memorizing the map” rather than “learning the compass”. This tendency results in brittle generalization under distribution shifts.
This observation motivates a different design principle: rather than learn an unconstrained update policy, the optimizer should learn an iterative dynamical system with an explicit stability-inducing bias. In response, we propose Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search through stability-inducing neural unrolling. Concretely, L2E models population evolution as an unrolled iterative solver [17, 49]. The inner loop follows a stability-biased update structure (e.g., averaged/residual updates with bounded transformations), while the outer loop meta-trains the operator to produce effective search trajectories. To balance global exploration with local refinement, L2E further introduces a gradient-derived composite solver. This solver adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner, enabling controlled progress under limited evaluations. From an architectural perspective, L2E parameterizes the learned evolutionary operator with a structured Mamba-based design to model population-wise interactions. By combining state-space modeling with lightweight routing over population statistics, the operator provides a scalable inductive bias for meta-learned evolutionary search. This stability-inducing formulation yields the behavioral contrast illustrated in Fig. 1. Specifically, while unconstrained learned heuristics may overfit training landscapes and exhibit unstable trajectories on unseen problems, L2E promotes controlled search dynamics by learning structured evolutionary updates via neural unrolling. This framework explicitly emphasizes stable dynamics, scalability to high-dimensional settings, and strong empirical generalization under distribution shifts. Our main contributions are:
  • Stability-Inducing Unrolling Formulation: We propose Learning to Evolve (L2E), a bilevel meta-optimization framework that learns evolutionary search via stability-inducing neural unrolling. The inner loop enforces structured population updates that are stability-oriented, while the outer loop meta-trains the operator to produce effective search trajectories across tasks.
  • Structured Operator and Composite Solver: We develop a Gradient-derived Evolutionary Composite (GEC) solver that adaptively fuses learned evolutionary proposals with proxy numerical guidance in a differentiable manner. We further introduce a structured Mamba operator to capture population-wise interactions, providing a scalable inductive bias in high-dimensional regimes.
  • Analysis and Empirical Validation: We analyze how the proposed unrolling structure induces controlled optimization dynamics, and validate L2E on synthetic benchmarks and real-world control tasks. Results show strong performance, scalability to high dimensions, and robust zero-shot generalization under distribution shifts.
@@
Figure 1: Conceptual contrast. Left: Unconstrained learned heuristics may overfit training landscapes, resulting in unstable trajectories on unseen problems. Right: L2E learns a stability-inducing unrolled evolutionary dynamics with bilevel meta-training, shaping the search trajectory toward controlled progress and improved robustness under distribution shifts.
@@ II-A Learning to Optimize
Learning to Optimize (L2O) combines the rigor of classical optimization with the capabilities of machine learning [3]. Specifically, rather than relying on hand-crafted heuristics, L2O seeks to parameterize update rules directly from data. Early works employed model-free architectures, such as RNNs [2] or MLPs [47], to map gradients to updates in a coordinate-wise manner. However, these methods often struggle with scalability and interpretability, primarily due to their opaque nature. In contrast, model-based L2O approaches embed structural priors into the learner. For instance, transformer-based optimizers [16] leverage self-attention to capture parameter dependencies. This methodology has further evolved into meta-learning frameworks that train optimizers to adapt across diverse tasks [45, 27]. Related directions include reinforcement learning controllers for scheduling [44] and gradient-guided mutation operators [32].
Despite these advances, existing L2O approaches primarily function as unconstrained policy networks. Specifically, they lack intrinsic geometric constraints. This deficiency often leads to unstable optimization dynamics and brittle generalization under distribution shifts. Consequently, these limitations highlight the need for structure-aware learned optimizers that promote controlled update behaviors.
@@ II-B Learnable Evolutionary Algorithms
Classical evolutionary algorithms and Bayesian Optimization methods have established the foundation for gradient-free optimization. Prominent examples include Evolutionary Strategies, such as CMA-ES [21], Differential Evolution [50], and L-SHADE [57]. Within the Bayesian domain, notable approaches include SAASBO [10], TuRBO [11], and HEBO [7]. These methods proved successful due to engineered covariance adaptation and population mechanisms [25]. However, their reliance on hand-crafted, non-differentiable rules limits their flexibility in high-dimensional or ill-conditioned landscapes. To enhance adaptability, recent research has progressively integrated learnability into evolutionary frameworks [71, 60, 29, 61, 28]. Initial efforts utilized surrogates, such as Gaussian Processes, to approximate objective landscapes [40]. Subsequent work has focused on embedding data-driven modules directly into the optimization loop. Examples include meta-evolved operators [41] and context-aware optimizers [44, 45]. Concurrently, researchers have employed Reinforcement Learning for dynamic operator selection [19, 18, 30]. Furthermore, Large Language Models are increasingly being explored for synthesizing heuristic code [55, 13, 23].
Despite this progress, learnable EAs face substantial challenges. Most methods optimize isolated components, rather than the unified evolutionary loop, leading to fragmented training and limited end-to-end coordination between gradient guidance and population search. As a result, robust zero-shot generalization and scalability to high-dimensional regimes remain difficult to achieve.
@@ II-C Neural Unrolling
Within the broader inverse problem community, Neural Unrolling (NU) has established itself as a principled methodology for bridging model-based iteration and deep learning. Fundamentally, NU reinterprets an iterative optimization method as a computational graph, wherein each layer corresponds to a single update step and algorithmic parameters function as trainable weights [69]. Although originally developed for signal and image processing, this paradigm has produced influential architectures, including LISTA for sparse coding [17], ADMM-Net for compressive sensing [68], and primal-dual networks for medical image reconstruction [1]. NU embeds domain-specific iterations into differentiable architectures. Recent advances have extended NU to complex visual perception tasks, such as image dehazing and rain removal, by incorporating optimization-inspired heuristics into network design to enhance feature interpretability [39, 37, 38]. Furthermore, in the realm of AutoML, differentiable approaches for Neural Architecture Search (NAS) implicitly leverage the unrolling concept to optimize structural parameters via bilevel gradients [34, 35, 49]. Consequently, it synergizes the interpretability of model-based methods with the adaptability of deep networks, thereby achieving robust performance across diverse problems.
Despite its prevalence in signal processing, NU remains underutilized within evolutionary computation. Although differentiable EAs exist, they typically rely on ad hoc relaxations rather than systematic unrollings of fixed-point iterations. Consequently, the field lacks a framework that unifies bilevel meta-learning with stability-oriented population dynamics in a principled and trainable manner.
@@ II-D Discussion
Existing methodologies navigate a trade-off between stability and adaptability. While classical EAs provide theoretical stability, they often lack the adaptability for navigating complex landscapes efficiently. Conversely, emerging L2O methods offer data-driven flexibility; however, they frequently exhibit instability and poor generalization to out-of-distribution landscapes, primarily due to the absence of intrinsic geometric constraints. To reconcile these diverging paradigms, we propose the L2E that reformulates the evolutionary loop as a neural unrolling process. This formulation integrates the expressivity of deep networks with controlled optimization dynamics. Consequently, it improves robustness under distribution shifts.
Figure 2: Overall framework of L2E. Left: L2E is formulated as a bilevel meta-optimization problem where the outer loop updates meta-parameters to shape the induced evolution dynamics. Middle: The evolutionary search is unrolled into a differentiable pipeline via AutoDiff to learn stable and effective search trajectories. Right: L2E implements a stability-inducing operator design that combines a learned evolutionary operator with numerical guidance, promoting controlled population updates.
@@ III Methodology
The overall architecture operates as a bilevel dynamical system, as visualized in Fig. 2. To instantiate this paradigm, the methodology comprises three core components. First, the Generalized Neural Unrolling (GNU) formulation establishes the theoretical foundation. Second, the Gradient-derived Evolutionary Composite (GEC) solver integrates gradient-based exploitation with population-based exploration. Finally, the structured Mamba neural operator efficiently parameterizes the optimization process in high-dimensional spaces, thereby improving computational scalability and promoting controlled optimization dynamics.
@@ III-A Generalized Neural Unrolling Paradigm
Conventional black-box optimizers often rely on heuristic mechanisms that can yield unstable dynamics on unseen landscapes. To promote controlled update behaviors, we reinterpret population evolution as a structured operator iteration, adopting a fixed-point inspired unrolling view based on the KM formulation [64]. Unlike methods that directly search for optima, the proposed approach optimizes the parameters of a learnable fixed-point operator. This operator implicitly governs the solution trajectory. Formally, let $f:\mathcal{X}\to\mathbb{R}$ denote the black-box objective function. We introduce a learnable non-expansive-inspired operator $\mathcal{NU}(\cdot;\bm{\omega})$, parameterized by $\bm{\omega}\in\bm{\Omega}$, whose fixed point approximates the minimizer of $f$. Consequently, we formulate this framework as a constrained bilevel optimization problem [36]:
@@
where $\ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega},\mathtt{Opt})$ is the meta-objective that evaluates the population state $\mathbf{x}$ induced by the optimizer parameter $\bm{\omega}$ under a prescribed evaluation protocol $\mathtt{Opt}$. Furthermore, $\text{Fix}(\mathcal{NU}):=\{\bm{\mathbf{x}}\mid\bm{\mathbf{x}}=\mathcal{NU}(\bm{\mathbf{x}};\bm{\omega})\}$ defines the set of fixed points. Here, $\bm{\omega}$ are outer-level meta-parameters that parameterize the optimizer, whereas $\mathbf{x}$ denotes the inner-level population state of candidate solutions.
To introduce a stability-biased update structure, we construct $\mathcal{NU}$ as an $\alpha$-averaged (KM-style) operator. Specifically, this formulation employs a convex combination of the identity and a learnable composite mapping $\mathcal{O}(\cdot,\bm{\omega})$, i.e.,
@@
The operator $\mathcal{NU}(\cdot;\bm{\omega})$ is defined as a composition of two components, modulated by an update relaxation parameter $\alpha\in(0,1]$. $\mathcal{O}_{\mathrm{evo}}$ is a learnable evolutionary proposal operator, and $\mathcal{O}_{\mathrm{num}}$ provides proxy numerical guidance (e.g., projection or smoothing) to regulate the update behavior. Together, the averaging step and the composite structure impose a stability-inducing bias on the unrolled trajectory, which empirically reduces the brittle dynamics often observed in unconstrained learned optimizers.